OpenStudy (anonymous):

help please? i will give medal and fan. Find line AB. round to the nearest tenth.

OpenStudy (imstuck):

There's a formula for that! In terms of what you have there, the formula would be \[(18)^{2}=x(x+12)\]with x as the line you are looking for. Do the math. 18*18=324. Distribute the x into the parenthesis and get x^2 + 12x. set this up as a polynomial and then factor it.\[x ^{2}+12x-324=0\]The answer I get using the quadratic formula is really wacky. It came out to be that x is\[-6+6\sqrt{10}\]which comes out to about 12.97 or 13. Are you sure you got the numbers on the lines correct?